An 8 kilogram bowling ball is rolling in a straight line toward you. If its momentum is 16 kg•m/sec, how fast it is traveling.


1
Expert's answer
2021-01-19T07:09:38-0500

By the definition of momentum, we have:


"p=mv,""v=\dfrac{p}{m}=\dfrac{16\ \dfrac{kgm}{s}}{8\ kg}=2\ \dfrac{m}{s}."

Answer:

"v=2\ \dfrac{m}{s}."